OpenStudy (anonymous):

find the equation of a circle with centre as (3,5) and radius as 8 units

OpenStudy (anonymous):

(x-3)^2+(y-5)^2=(8)^2 \[x ^{2}+y ^{2}-6x-10y-30=0\]

OpenStudy (anonymous):

(x-3)\[\left( x-3 \right)^{2} +\left( y-5 \right)^{2} =8^{2}\]

OpenStudy (anonymous):

u guys r making me confused ! which is true ?

OpenStudy (anonymous):

\[\left( x-3 \right)^{2}+\left( y-5 \right)^{2}= 8^{2}\]